# The Procrustean Bed of Time Series: The Optimization Bias of Point-wise Loss

### v1

@@ Abstract
Optimizing time series models via point-wise loss functions (e.g., MSE) relying on a flawed point-wise independent and identically distributed (i.i.d.) assumption that disregards the causal temporal structure, an issue with growing awareness yet lacking formal theoretical grounding. Focusing on the core independence issue under covariance stationarity, this paper aims to provide a first-principles analysis of the Expectation of Optimization Bias (EOB), formalizing it information-theoretically as the discrepancy between the true joint distribution and its flawed i.i.d. counterpart. Our analysis reveals a fundamental paradigm paradox: the more deterministic and structured the time series, the more severe the bias by point-wise loss function. We derive the first closed-form quantification for the non-deterministic EOB across linear and non-linear systems, and prove EOB is an intrinsic data property, governed exclusively by sequence length and our proposed Structural Signal-to-Noise Ratio (SSNR). This theoretical diagnosis motivates our principled debiasing program that eliminates the bias through sequence length reduction and structural orthogonalization. We present a concrete solution that simultaneously achieves both principles via DFT or DWT. Furthermore, a novel harmonized $\ell_{p}$ norm framework is proposed to rectify gradient pathologies of high-variance series. Extensive experiments validate EOB Theory’s generality and the superior performance of debiasing program.
@@
Figure 1: Motivation of our work. The use of a point-wise loss function leads to the deep learning model to presume an independent and identical distribution as an approximation for the true joint distribution of data. It is biased.
@@ 1 Introduction
Time series analysis is a foundational task in machine learning (Zhang et al., 2024), with applications spanning economics (Zhang et al., 2025; Bao et al., 2025), industry (Cai et al., 2024), and healthcare (Li et al., 2024). The performance of deep learning models in this domain primarily depends on two core perspectives: the expressive power of the network architecture and the rationality of the training paradigm, particularly the loss function design.
@@
Model expressiveness has been the primary focus of the research community. This has led to a vibrant and rapidly evolving landscape of complex architectures, from foundational RNNs (Kong et al., 2025), CNNs (Donghao & Xue, 2024) and GNNs (Chen et al., 2024b) to the current struggle for supremacy among the Transformer family (Liu et al., 2024; Chen et al., 2024a; Jin et al., 2024b), advanced MLPs (Ekambaram et al., 2023) and Diffusion (Rasul et al., 2021) in temporal domain. Concurrently, modeling in the frequency domain has emerged as a powerful new force (Yi et al., 2023b, 2025; Yang et al., 2024; Zhong et al., 2025; Cheng et al., 2025; Xu et al., 2024; Crabbé et al., 2024), with lightweight linear models demonstrating surprisingly superior performance over heavyweight architectures.
@@
In stark contrast, the rationality of training paradigm that arguably determines the upper limit of model performance, has remained largely unexamined. The optimization of models via point-wise loss functions (e.g., MSE, MAE) in temporal domain is the standard paradigm for almost forecasting, imputation, and anomaly detection tasks (Jin et al., 2024a). While intuitive, this approach is built on a heuristic point-wise independent and identically distributed (i.i.d.) assumption, which fundamentally conflicts with the causal, temporal structure that defines time series data with dynamic characteristics, as illustrated in Figure 1. This “i.i.d.” assumption contains two distinct flaws: the identically distributed assumption, which is violated by dynamic non-stationarity, and the independence assumption, which is violated by temporal causality. We argues that the independence issue is the more foundational one, as it has a fundamental conflict with the contextual causality inherent in time series. We therefore focus on isolating this core independence issue, which allows us to build a rigorous theory under the well-defined conditions of covariance-stationarity.
Recent studies have begun to confront this issue. Works such as FreDF (Wang et al., 2025b) and Time-o1 (Wang et al., 2025a) explore this issue from the perspective of “label autocorrelation”, yet their analysis remains preliminary; the bias is defined as a posteriori effect dependent on a known model output, failing to identify the core, a prior factors that govern it. Another line of research engineers more sophisticated loss functions. For instance, PSLoss (Kudrat et al., 2025) incorporates local similarity measures akin to PatchTST (Nie et al., 2023), while DBLoss (Qiu et al., 2025) explicitly decomposes trend and periodic components to balance their optimization contributions. Despite their empirical success, liking the faint light before dawn, the motivation for these methods remains largely heuristic, lacking a systematic theoretical framework to explain the underlying optimization pathology.
@@
To cut through this fog of heuristic exploration, this paper ambitiously establishes the first-principled, systematic theoretical analysis of the optimization bias stemming from the point-wise independence assumption. Wielding the tools of information theory and stochastic process modeling (Ross, 2023), we chart a clear path from diagnosis to prescription. We begin by defining the Expectation of Optimization Bias (EOB) as the probabilistic divergence, i.e., KL-divergence, between the true joint distribution and the flawed factorized distribution assumed by point-wise loss function. This formalization allows us to uncover the bounds of EOB and reveal a fundamental Paradigm Paradox: the more deterministic the process, the more severely it is penalized by point-wise loss, a finding that depends purely on the data’s intrinsic properties.
@@
Our analysis proves that the lower bound of the EOB is governed by the non-deterministic component of the process. We therefore proceed to quantify this lower bound for both linear (parametric autoregressive model and non-parametric multivariate Gaussian model) and non-linear (Gaussian mixture model) systems, abstracting its decisive factors (i.e., sequence length and structural signal-to-noise ratio). Based on this solid theoretical diagnosis, we propose a Principled Debiasing Program that not only provides a unifying framework for existing methods but also motivates a novel and robust solution. Concretely, we instantiate our program through Discrete Fourier Transform (DFT). Moreover, we improve the optimization robustness in frequency domain via a novel harmonized $\ell_{p}$ norm framework to resolve the gradient pathologies.
To summarize, our contributions include:
@@
  • Formalization of Paradigm Paradox: This work is the first to formalize systematic analysis of the optimization bias inherent to point-wise loss functions, which stems from their heuristic i.i.d. assumption. We derive bounds on EOB, revealing a fundamental paradigm paradox: the more deterministic a time series, the more severely it is penalized by the point-wise loss.
@@
  • Quantification of EOB Lower Bound: We present the first closed-form quantification of non-deterministic EOB for linear process, unifying the analysis of parametric and non-parametric models. We further extend this theory to non-linear processes.
@@
  • Principled Debiasing Program: We propose an essential principle that eliminates the bias. Furthermore, a concrete solution is introduced using DFT or DWT with a harmonized $\ell_{p}$ norm optimization to resolve gradient pathologies of high-variance series.
  • Comprehensive Empirical Validation: Experiments on synthetic and real-world benchmarks validates the generality of our EOB Theory and debiasing program.
@@ 2 Optimization Bias of Point-wise Loss
Our theory first establishes a formal definition and the lower bound for EOB in Section 2.1. We then quantify the non-deterministic EOB in linear systems through both dynamic and correlational insights, using a parametric autoregressive model (Section 2.2.1) and a non-parametric multivariate Gaussian model (Section 2.2.2), and derive a unified expression. Finally, we extend these linear findings to non-linear conditions via Gaussian mixture model in Section 2.3.
2.1 Formalization of Probability Distribution Divergency
In machine learning, the minimization of a training loss is implicitly equivalent to maximizing the likelihood of a probabilistic model, $q(x_{1:T})$, which serves as an approximation of the true data-generating distribution, $p(x_{1:T})$.The ubiquitous use of point-wise loss functions (e.g, MSE, MAE) introduces a critical flaw by incorrectly assuming that each data point in a sequence is conditionally independent. We term this the point-wise independence assumption:
@@ 2.1 Formalization of Probability Distribution Divergency
However, the defining characteristic of a time series is its temporal structure, where observations are sequentially dependent. The true joint probability distribution, $p(x_{1:T})$, must therefore be expressed using the chain rule of probability to respect this inherent auto-correlation:
@@
This reveals a fundamental discrepancy between the training objective and the true data structure. By decomposing the true log-likelihood, $\log p(x_{1:T})$, we can precisely quantify the divergence introduced by the point-wise independence assumption:
@@
We define this correction term as the Optimization Bias $\mathcal{B}$ of point-wise loss. It represents the accumulated information gain from conditioning on the past, a factor completely overlooked when a model is trained by minimizing a simple sum of independent errors.
To analyze the properties of this bias, we can bound its expectation by considering the composition of the time series.
@@
(Bounds on the Expectation of Optimization Bias) Let $\{x_{t}\}$ be a process admitting a Cramér decomposition $x_{t}=v_{t}+z_{t}$, where $\{v_{t}\}$ is the deterministic component and $\{z_{t}\}$ is the purely stochastic component. Assume the following conditions hold:
1. Determination: $v_{t}$ is perfectly predictable from its past $v_{1:t-1}$, such that conditional entropy $H(v_{t}|v_{1:t-1})=$-$\infty$.
2. Independence: $\{v_{t}\}$ and $\{z_{t}\}$ are independent.
3. Information Equivalence: The information contained in the history $x_{1:t-1}$ is equivalent to that in $\{v_{1:t-1},z_{1:t-1}\}$.
Then, the expectation of optimization bias of the sequence $x_{t}$ over a duration $T$, $\mathbb{E}[\mathcal{B}]$, is bounded as follows:
$\mathbb{E}\left[\sum_{t=2}^{T}\log\frac{p(z_{t}|z_{1:t-1})}{p(z_{t})}\right]\leq\mathbb{E}[\mathcal{B}]\leq\mathbb{E}\left[\sum_{t=2}^{T}\log\frac{p(v_{t}|v_{1:t-1})}{p(v_{t})}\right]$ (4)
@@
The upper bound in Theorem 2.1 (Cramér, 1946), driven by the deterministic component’s bias, approaches infinity for any non-trivial deterministic process. This leads to a counterintuitive conclusion, formalized as a paradox.
@@ 2.2 Linear Modeling of Non-deterministic EOB
2.2.1 Parametric Autoregressive Model
@@
(Quantification of Non-deterministic EOB via $AR(p)$) For any covariance-stationary Gaussian process driven by $AR(p)$, the non-deterministic EOB, $\mathbb{E}[\mathcal{B}_{z}]$, incurred by a point-wise loss function is quantified by: $\mathbb{E}[\mathcal{B}_{z}]=\frac{T-p}{2}\log\frac{\sigma_{z}^{2}}{\sigma_{\epsilon}^{2}}=\frac{T-p}{2}\log\frac{1}{1-\sum_{i=1}^{p}\phi_{i}\rho_{i}}$ (7) where $\phi_{i}$ and $\rho_{i}$ are the autoregressive coefficient and autocorrelation function at lag $i$, respectively.
@@
This result is profound and can be deconstructed into two key factors: Sequence Length (TpT-p), over which the EOB accumulates linearly, and variance ratio ($\sigma_{z}^{2}/\sigma_{\epsilon}^{2}$). The latter term is a significant statistics, formalized as follows:
@@
(Structural Signal-to-Noise Ratio) For any covariance-stationary process, SSNR is the ratio of the total unconditional variance ($\sigma_{z}^{2}$) to its one-step-ahead optimal prediction error variance ($\sigma_{\epsilon}^{2}$): $\textit{SSNR}=\frac{\sigma_{z}^{2}}{\sigma_{\epsilon}^{2}}=\frac{\operatorname{Var}(z_{t})}{\operatorname{Var}(z_{t}-\mathbb{E}[z_{t}|z_{1:t-1}])}$ (8) where $\mathbb{E}[z_{t}|z_{1:t-1}]$ is the optimal prediction of $z_{t}$ given all past information, and the denominator represents the variance of the process’s irreducible innovation.
Conceptually, SSNR measures the ratio of a process’s total uncertainty to its residual uncertainty. A high SSNR signifies a process with strong internal structure and high predictability. A key property of SSNR within the $AR(p)$ model is its exclusive dependence on the intrinsic autoregressive coefficients $\phi_{i}$, rendering it a pure measure of the series’s internal structure, independent of the noise scale. The relationship between our defined SSNR and classical SNR is further detailed in Appendix C.
2.2.2 Non-Parametric Multivariate Gaussian Model
While the preceding analysis provides a closed-form solution for a parametric $AR(p)$ process, we now generalize this result to a broader, non-parametric condition. Considering any linear, covariance-stationary process can be driven by a Multivariate Gaussian Model (MGM), it leads to an elegant and fundamental principle governing $\mathbb{E}[\mathcal{B}_{z}]$.
$\mathrm{z}\sim\mathcal{N}(\boldsymbol{\mu},\Sigma).$ (9)
@@ 2.2.2 Non-Parametric Multivariate Gaussian Model
(Quantification of Non-deterministic EOB via MGM) For any stochastic, covariance-stationary process modeled by a multivariate Gaussian distribution, the non-deterministic EOB, is determined solely by the determinant of correlation matrix, $|R|$: $\mathbb{E}[\mathcal{B}_{z}]=-\frac{1}{2}\log|R|.$ (10)
@@
This result provides a profound insight, unifying the parametric and non-parametric perspectives. It reveals that the total non-deterministic EOB is composed of two parts: a primary term, $\frac{T-p}{2}\log(\text{SSNR})$, which represents the bias accumulated during the TpT-p steady-state steps, and a constant correction term, $\frac{1}{2}\log|R_{p}|$, accounting for the bias from the initial $p$-step transient phase. This unification culminates in a central theorem for quantifying the EOB.
@@
(Quantification of Non-deterministic EOB) For any covariance-stationary process, the non-deterministic EOB, $\mathbb{E}[\mathcal{B}_{z}]$, incurred by a point-wise loss is fundamentally determined by the sequence length $T$ and its SSNR, and can be expressed as: $\mathbb{E}[\mathcal{B}_{z}]=\frac{T}{2}\log(\textit{SSNR})+c$ (15) where $c=-\frac{1}{2}\log(|R_{p}|\cdot(\textit{SSNR})^{p})$ is a constant only relying on the initial short-term memory of the process.

### v2

@@ Abstract
Optimizing time series models via point-wise loss functions (e.g., MSE) relying on a heuristic point-wise i.i.d. assumption disregards the causal temporal structure. Focusing on the core independence issue under covariance stationarity, this paper aims to provide a first-principles analysis of the Expectation of Optimization Bias (EOB). Our analysis reveals a fundamental paradigm paradox: The more deterministic and structured the time series, the more severe the bias incurred by point-wise loss function. We derive the first closed-form quantification for the non-deterministic EOB across linear and non-linear systems, and prove EOB is an intrinsic data property, governed exclusively by sequence length and the defined Structural Signal-to-Noise Ratio. This theoretical discovery motivates our principled debiasing program that eliminates the bias through sequence length reduction and structural orthogonalization. We present a concrete solution via DFT or DWT, and propose a novel harmonized $\ell_{p}$ norm framework to rectify gradient optimization pathologies of high-variance sequences. Extensive experiments validate EOB Theory’s generality and the superior performance of debiasing program, achieving 5.2% and 5.1% average improvement of MSE and MAE conducted on the iTransformer across 11 datasets, respectively. [Code Link]
@@
Figure 1: Motivation of our work. Point-wise losses implicitly treat each time step as i.i.d., approximating the true joint distribution with a fully factorized surrogate. This mismatch induces a systematic optimization bias.
@@ 1 Introduction
Time series analysis is a foundational pillar in machine learning (Zhang et al., 2024), with applications spanning economics (Zhang et al., 2025; Bao et al., 2025), industry (Cai et al., 2024), and healthcare (Li et al., 2024). The performance of deep learning models in this domain primarily hinges on two core perspectives: the expressive power of the network architecture and the rationality of the training paradigm, particularly the loss function design.
@@
Model expressiveness has been the primary focus of the research community. This has led to a vibrant and rapidly evolving landscape of complex architectures, from foundational RNNs (Kong et al., 2025), CNNs (Luo & Wang, 2024) and GNNs (Chen et al., 2024b) to the current struggle for supremacy among the Transformer family (Liu et al., 2024; Chen et al., 2024a; Jin et al., 2024b), advanced MLPs (Ekambaram et al., 2023) and Diffusion (Rasul et al., 2021) within the temporal domain. Concurrently, modeling in the frequency domain has emerged as a powerful new force (Yi et al., 2023b, 2025; Yang et al., 2024; Zhong et al., 2025; Cheng et al., 2025; Xu et al., 2024; Crabbé et al., 2024), with lightweight linear models demonstrating surprisingly superior performance relative to heavyweight architectures.
@@
In stark contrast, the rationality of the training paradigm that arguably determines the upper limit of model performance, has remained largely unexamined. The optimization of models via point-wise loss functions (e.g., MSE, MAE) in the temporal domain is the standard paradigm for nearly all forecasting, imputation, and anomaly detection tasks (Jin et al., 2024a). While intuitive, this approach is built on a heuristic point-wise independent and identically distributed (i.i.d.) assumption, which fundamentally conflicts with the causal, temporal structure that defines time series data with dynamic characteristics, as illustrated in Figure 1. This assumption entails two obvious flaws: the identically distributed assumption, which is violated by dynamic non-stationarity, and the independence assumption, which is violated by temporal causality. We argue that the independence issue is the more foundational one, as it has a fundamental conflict with the contextual causality inherent in time series. We therefore focus on isolating this core independence issue, enabling the development of a rigorous theory under the well-defined conditions of covariance-stationarity.
Recent literature has initiated preliminary investigations into these optimization biases. Specifically, FreDF (Wang et al., 2025b) and Time-o1 (Wang et al., 2025a) characterize this issue via “label autocorrelation”, yet their perspective is limited to a posteriori observation of model outputs, leaving the root a priori causes unidentified. Parallel efforts, including PSLoss (Kudrat et al., 2025) and DBLoss (Qiu et al., 2025), introduce empirical statistical constraints to regularize training. Despite their success, the motivation behind these loss functions remains largely intuitive rather than principled. They capture the “faint light” of improved performance but fail to provide a systematic theoretical scaffolding to explain the underlying mechanics of objective-induced bias. There remains a critical gap between empirical correction and theoretical understanding of temporal optimization pathologies.
@@
To move beyond this fog of heuristic exploration, this work ambitiously establishes the first-principled, systematic theoretical analysis of the optimization bias stemming from the point-wise independence assumption. Wielding the tools of information theory and stochastic process modeling (Ross, 2023), we chart a clear path from diagnosis to prescription. We begin by defining the Expectation of Optimization Bias (EOB) as the probabilistic divergence, i.e., KL-divergence, between the true joint distribution and the flawed factorized distribution implicitly assumed by point-wise loss functions. This formalization allows us to derive the bounds of EOB and reveal a fundamental paradigm paradox: The more deterministic the process is, the more severely it is penalized by point-wise loss, a finding that governed strictly by the data’s intrinsic properties.
@@
Our analysis proves that the lower bound of the EOB is governed by the non-deterministic component of the process. Accordingly, we quantify this lower bound for both linear (parametric autoregressive model and non-parametric multivariate Gaussian model) and non-linear (Gaussian mixture model) systems, identifying its decisive determinants (i.e., sequence length and structural signal-to-noise ratio). Based on this solid theoretical diagnosis, we propose a Principled Debiasing Program that not only provides a unifying framework for existing methods while facilitating a novel and robust solution. Concretely, we instantiate this program using DFT and DWT. Moreover, we improve the optimization robustness for high-variance sequences via a novel harmonized $\ell_{p}$ norm framework to mitigate the gradient pathologies.
To summarize, our theoretical contributions include:
@@
  • Formalization of Paradigm Paradox: This work provides the first systematic formalization of the optimization bias inherent in point-wise loss functions, tracing its roots to the heuristic i.i.d. assumption. We derive the bounds of EOB, revealing a fundamental paradigm paradox: the more deterministic a time series is, the more severely it is penalized by the point-wise loss.
@@
  • Quantification of EOB Lower Bound: We present the first closed-form quantification of non-deterministic EOB for linear Gaussian processes, unifying the analysis of parametric and non-parametric models. We further extend this theory to non-linear processes.
@@
  • Principled Debiasing Program and Verification: We propose an essential principle to eliminate the bias. Specifically, our implementation using DFT or DWT achieves average reductions of 5.2% in MSE and 5.1% in MAE across 11 datasets when deployed on iTransformer. Moreover, we propose a novel harmonized $\ell_{p}$ norm for robust optimization. Crucially, the optimization effect is independent of architecture and dataset.
@@ 2 Optimization Bias of Point-wise Loss
Our theory first establishes the formal definition and lower bound of EOB in Section 2.1. We then quantify the non-deterministic EOB in linear systems through both dynamic and correlational insights, using a parametric autoregressive model (Section 2.2.1) and a non-parametric multivariate Gaussian model (Section 2.2.2), and derive a unified expression. Finally, we extend these linear findings to non-linear conditions via Gaussian mixture model in Section 2.3.
2.1 Formalizing Optimization Bias via Distribution Divergence
In machine learning, the minimization of the training loss is implicitly equivalent to maximizing the likelihood under a probabilistic model, $q(x_{1:T})$, which serves as an approximation of the true data-generating distribution, $p(x_{1:T})$. The ubiquitous use of point-wise loss functions (e.g., MSE, MAE) introduces a critical flaw by incorrectly assuming that each data point in a sequence is conditionally independent. We term this the point-wise independence assumption:
@@ 2.1 Formalizing Optimization Bias via Distribution Divergence
However, the defining characteristic of a time series is its temporal structure, where observations are sequentially dependent. The true joint probability distribution, $p(x_{1:T})$, must therefore be expressed using the chain rule of probability to respect this inherent autocorrelation:
@@
This reveals a fundamental discrepancy between the training objective and the true data structure. By decomposing the true log-likelihood, $\log p(x_{1:T})$, we quantify the divergence introduced by the point-wise independence assumption:
@@
We define this correction term as the Optimization Bias $\mathcal{B}$ of the point-wise loss. It represents the accumulated information gain from conditioning on the past, a factor completely overlooked when a model is trained by minimizing a simple sum of independent errors.
To analyze the properties of this bias, we bound its expectation by considering the structural composition of the series.
@@
(Bounds on the Expectation of Optimization Bias) Let $\{x_{t}\}$ be a discrete-time stochastic process of continuous random variables admitting a Cramér decomposition $x_{t}=v_{t}+z_{t}$, where $\{v_{t}\}$ is the deterministic component and $\{z_{t}\}$ is the purely stochastic component. Assume the following conditions hold: A.1 Continuous Determinism, A.2 Independence, and A.3 Structural Identifiablity.
Then, the expectation of optimization bias for the sequence $x_{t}$ over duration $T$, denoted as $\mathbb{E}[\mathcal{B}]$, is bounded by the mutual information of its components:
$\underbrace{\mathbb{E}\left[\sum_{t=2}^{T}\log\frac{p(z_{t}|z_{1:t-1})}{p(z_{t})}\right]}_{\text{Lower Bound: }I(z_{1:T})}\leq\mathbb{E}[\mathcal{B}]\leq\underbrace{\mathbb{E}\left[\sum_{t=2}^{T}\log\frac{p(v_{t}|v_{1:t-1})}{p(v_{t})}\right]}_{\text{Upper Bound: }I(v_{1:T})\to\infty}$ (4)
@@
The upper bound in Theorem 2.1 (Cramér, 1946), driven by the bias of deterministic component, approaches infinity for any non-trivial deterministic process. This leads to a counterintuitive conclusion, formalized as a paradox.
@@ 2.2 Linear Modeling of Non-deterministic EOB
2.2.1 Param. Autoregressive Model
@@
(Quantification of Non-deterministic EOB via $AR(p)$) For any covariance-stationary Gaussian process driven by $AR(p)$, the non-deterministic EOB, $\mathbb{E}[\mathcal{B}_{z}]$, incurred by a point-wise loss is given as: $\mathbb{E}[\mathcal{B}_{z}]=\frac{T-p}{2}\log\frac{\sigma_{z}^{2}}{\sigma_{\epsilon}^{2}}+C_{p}$ (7) where $C_{p}$ is a constant about intrinsic data property.
@@
This result is profound and can be deconstructed into two key factors: Sequence Length (TpT-p), over which the EOB accumulates linearly, and variance ratio ($\sigma_{z}^{2}/\sigma_{\epsilon}^{2}$). The latter term is a pivotal statistic, formalized as follows:
@@
(Structural Signal-to-Noise Ratio (SSNR)) For any process, SSNR is the ratio of total unconditional variance ($\sigma_{z}^{2}$) to its one-step-ahead optimal prediction error variance ($\sigma_{\epsilon}^{2}$): $\textit{SSNR}=\frac{\sigma_{z}^{2}}{\sigma_{\epsilon}^{2}}=\frac{\operatorname{Var}(z_{t})}{\operatorname{Var}(z_{t}-\mathbb{E}[z_{t}|z_{1:t-1}])}.$ (8)
Specifically, $\mathbb{E}[z_{t}|z_{1:t-1}]$ is the optimal prediction of $z_{t}$ given all past information, and the denominator represents the variance of the process’s irreducible innovation.
Conceptually, the SSNR measures the ratio of a process’s total uncertainty to its residual uncertainty. A high SSNR signifies a process with strong internal structure and high predictability. A key property of SSNR within the $AR(p)$ model is its exclusive dependence on the intrinsic autoregressive coefficients $\phi_{i}$, rendering it a pure measure of the series’ internal structure, independent of the noise scale. The relationship between our defined SSNR and the classical SNR is further detailed in Appendix C.
2.2.2 Non-Param. Multivariate Gaussian Model
While the preceding analysis provides a closed-form solution for a parametric $AR(p)$ process, we now generalize this result to a broader, non-parametric setting. Since any linear, covariance-stationary process can be modeled as a Multivariate Gaussian Model (MGM), it leads to an elegant and fundamental principle governing $\mathbb{E}[\mathcal{B}_{z}]$.
$\boldsymbol{\mathrm{z}}\sim\mathcal{N}(\boldsymbol{\mu},\Sigma).$ (9)
@@
(Quantification of Non-deterministic EOB via MGM) For any covariance-stationary stochastic process governed by a multivariate Gaussian model, the non-deterministic EOB, $\mathbb{E}[\mathcal{B}_{z}]$, is determined solely by the determinant of correlation matrix, $|R|$: $\mathbb{E}[\mathcal{B}_{z}]=-\frac{1}{2}\log|R|.$ (10)
@@
This result provides a profound insight, unifying the parametric and non-parametric perspectives. It reveals that the total non-deterministic EOB is composed of two parts: a primary term, $\frac{T-p}{2}\log(\textit{SSNR})$, which represents the bias accumulated during the TpT-p steady-state steps, and a constant correction term, $\frac{1}{2}\log|R_{p}|$, accounting for the bias from the initial $p$-step transient phase. This unification culminates in a central theorem for quantifying the EOB.
@@
(Quantification of Non-deterministic EOB) For any Gaussian and covariance-stationary stochastic process, the non-deterministic EOB, $\mathbb{E}[\mathcal{B}_{z}]$, incurred by a point-wise loss is fundamentally determined by the sequence length $T$ and its SSNR, quantified as: $\mathbb{E}[\mathcal{B}_{z}]=\frac{T}{2}\log(\textit{SSNR})+c$ (15) where $c=-\frac{1}{2}\log(|R_{p}|\cdot(\textit{SSNR})^{p})$ is a constant only relying on the initial short-term memory of the process.